Dividing Partnership Net Income

Required:

Steve Queen and Chelsy Boxer formed a partnership, dividing income as follows:

  1. Annual salary allowance to Queen of $95,700.
  2. Interest of 7% on each partner's capital balance on January 1.
  3. Any remaining net income divided to Queen and Boxer, 1:2.

Queen and Boxer had $48,000 and $111,000, respectively, in their January 1 capital balances. Net income for the year was $165,000. How much is distributed to Queen and Boxer?

Note: Compute partnership share.
Queen: $
Boxer: $
Answer #1

The partnership income is to be divided as per the partnership agreement on the basis of specified terms and conditions. To distribute net income to the partners, interest on loan, salary and allowance, commission to partners along with interest on capital shall be deducted from net profit to arrive at distributable income. Such income arrived is distributed in the agreed ratio of partners.

Particulars

Queen

Boxer

Total

Net Profit earned

         165,000

Less : Salary allowance

         95,700

         (95,700)

Remaining profits

           69,300

Less: Interest on capital (A)

           3,360

           7,970

         (11,130)

Balance profits

           58,170

Share of profits 1:2 (B)

         19,390

         38,780

         (58,170)

Total share of income of partners (A + B)

         22,750

         46,750

                    -  

Computation of interest on capital:

Queen = 48,000 * 7% = $ 3,360

Boxer = 111,000 * 7% = $ 7,970

kindly upvote